# Spectral Mode Enhancement in Coherent-harmonic Dual-comb Spectroscopy Enables Exceeding 300-fold averaging-time reduction

### v3

@@ Abstract
Dual-comb spectroscopy (DCS) is a novel Fourier-transform spectroscopy not relying on mechanical scanning and capable of simultaneously achieving high speed, high spectral resolution, and broad optical bandwidth. Nevertheless, it suffers from low signal-to-noise ratio (SNR) per single acquisition due to the dynamic range limitation of photodetectors imposed by the high-peak-power mode-locked pulses, making coherent averaging an essential means to improve SNR, at the price of compromising the exceptional time resolution and placing more stringent demands on mutual coherence and stability. In this study, a novel approach to enhance SNR by exploiting the spectral mode enhancement mechanism in coherent-harmonic pulses is demonstrated. As a proof-of-concept, two frequency combs with mode spacing of $\sim$12.5 MHz, operated at a 20th harmonic repetition rate of $\sim$250 MHz, are employed, demonstrating a $>$300-fold reduction in averaging time for comparable SNR in conventional DCS. This reduction is expected to be further enhanced through integration with ultra-high repetition rate combs like microresonator combs. This new approach promises both a recovery of the inherent high-speed capability and a mitigation of the coherence-time requirements, thereby making it possible to significantly facilitate subsequent DCS investigations and field deployments.
@@ I Introduction
Dual-comb spectroscopy (DCS) is a remarkable novel Fourier-transform spectroscopy technique powered by optical frequency comb (OFC) 1. Over the past decade, it has been increasingly recognized as an important spectroscopic tool in fields that require high speed, broad bandwidth, high frequency resolution, and high frequency accuracy, such as in investigations of physical turbulent environments 2, 3, 4, 5, 6 and chemical and biological dynamic processes 7, 8, 9, 10. By employing two coherent OFCs with a slight repetition rate difference ($\Delta f_{\mathrm{rep}}$), and inherently free from mechanical moving parts, DCS can capture an entire interferogram (IGM), yielding a spectrum with a full spectral resolution set by the repetition rate of the combs. The time frame for this capture is as short as $1/\Delta f_{\mathrm{rep}}$, and with large $\Delta f_{\mathrm{rep}}$ configurations, this leads to theoretical temporal resolution reaching microsecond or even nanosecond timescales 11, 12, 13.
@@
However, in typical practical DCS systems, the high peak power of mode-locked pulses often leads to photodetector saturation or nonlinear response, thus restricting the average power of the combs to a low level. Consequently, given the vast number of comb teeth inherent to broadband spectra, the power per comb tooth is typically limited to the nanowatt range 14, 15, ultimately resulting in extremely low SNR per single acquisition, which necessitates coherent averaging or multi-period acquisition to enhance the SNR 16, 15, 17, 12. Unfortunately, the SNR gain achieved through averaging scales proportionally to the square root of the number of averages. For instance, to improve the SNR by a factor of 10, a hundred averages are needed. Obviously, this process not only significantly compromises the exceptional time-resolution of DCS by several orders of magnitude, but also places greater demands on the long-term mutual coherence and stability of the system, which greatly increases the implementation complexity and cost.
@@
To mitigate this issue, one approach involves employing sequential spectrum acquisition using multiple photodetectors 14 or averaging spectral copies instead of IGMs 18. However, the SNR enhancement offered by these two methods has been limited by experimental complexity or sampling bandwidth constraints, demonstrating averaging-time reduction of less than one order of magnitude. Another early approach is to stretch the pulses of the probe comb using a chirped fiber Bragg grating 19. This method can effectively reduce the dynamic range of an IGM; however, it has limited impact on the nonlinearity or saturation induced by pulses, as the detector’s response time significantly exceeds the pulse duration. More recently, the utilization of $\sim$ 1-GHz-repetition-rate combs 12, 20, 13, 21 offers higher average power and larger $\Delta f_{\mathrm{rep}}$. Nevertheless, a $\sim$ 1-GHz repetition rate had struck a delicate balance between measurement speed and spectral resolution 22. A higher repetition rate inherently implies a lower spectral resolution, which limits further increases in repetition rate.
From this perspective, employing frequency combs analogous to certain harmonic mode-locked lasers 23, whose pulse repetition rate ($f_{\mathrm{rep}}$) operates on the harmonics of the laser cavity’s fundamental repetition rate, thus possibly featuring a high $f_{\mathrm{rep}}$ but a low mode spacing ($f_{\mathrm{ms}}$), may offer a promising avenue. On the one hand, this configuration can enable higher average power within the same dynamic range. On the other hand, the phenomenon that the pulse repetition rate exceeding the mode spacing likely points to an underlying mechanism of the harmonic pulses coherently enhancing their common spectral modes. If this mechanism can be effectively transferred to interferograms in DCS, i.e., generate multiple centerbursts within the time period of $1/\Delta f_{rep}$ to coherently enhance RF comb modes, it would be possible to achieve the desired SNR gain, thereby potentially refraining from extended IGMs coherent averaging.
Here we find pulse temporal phase modulation (TPM) using electro-optical phase modulator (EOPM) provides an ideal means to realize such combs unaffected by supermode noise 32. This TPM has been employed for the comb densification technique 25, 26, 27, 28, 29, 30, 31 based on spectral self-imaging effect 24 over the past decade or so, which involves applying periodic TPM to the pulses, inducing evenly spaced ”self images” of the original comb lines on the spectrum, thereby densifying the comb. In 2016, N.B. Hébert et al. 33 first introduced this densification technique in DCS and presented an enhanced sub-MHz spectral resolution from a 100-MHz comb, using a pseudo-random binary sequence (PRBS) modulation. In 2024, C. Q.-Galan et al. 34 extended this method to 1-GHz gain-switched OFCs, achieving $\sim$ 3.9 MHz resolution. Despite this, both of these two studies primarily focused on the spectral resolution enhancement, merely treating such combs as conventional OFCs with small mode spacing. The underlying harmonic-enhancement mechanism and the descriptions of such special dual-comb have not been elucidated.
In this work, we derived a coherent-harmonic-pulse superposition model in the frequency domain to elucidate this mechanism, and presented the time-domain and equivalent frequency-domain descriptions of two architectures for this special DCS, with a focus on their respective SNR gains. To align with its intuitive physical picture and distinguish it from conventional dual-comb spectroscopy (C-DCS), we term this approach coherent-harmonic-enhanced dual-comb spectroscopy (CH-EDCS). As a proof of concept, we perform transmission spectrum measurements of a hydrogen cyanide ($\mathrm{H^{13}C^{14}N}$) gas cell the two proposed architectures and, for comparison, a C-DCS architecture. The experiments were based on two commercial 250 MHz combs, which were configured using TPM and a pulse-picking technique 41 to operate at a spectral resolution of 12.5 MHz. As a result, we achieved an SNR gain of up to $\sim$ 18 times per single acquisition, equivalent to a $>$300-fold reduction in averaging time, demonstrating the significant potential for SNR enhancement of this approach.
Fig. 1: Concepts of conventional DCS (C-DCS) and C-H-DCS. (a) Time-domain description of C-DCS. The pulse train (red) of comb 1 progressively steps across another pulse train (blue) with a slight repetition rate difference $\Delta f_{\mathrm{s}}$, generating an interferogram with a single centerburst in a period of $1/\Delta f_{\mathrm{s}}$. When the temporal interval between the two pulses is large, effective interference does not occur.(b) The equivalent frequency-domain description of the C-DCS. The beating of two combs produces an RF comb in the radio-frequency domain. (c) Time-domain description of the C-H-DCS. The pulse of the conventional optical frequency comb (C-OFC) with a pulse repetition rate of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$, steps across the four pulses of the coherent-harmonic optical frequency comb (H-OFC), which have phases of $[\pi,0,0,0]$ and a pulse repetition rate of $4f_{\mathrm{ms}}$. This generates four centerbursts within an interferogram period of $1/\Delta f_{\mathrm{s}}$, which inherit the $[\pi,0,0,0]$ phases. (d) The equivalent frequency-domain description of the C-H-DCS. Two combs with mode spacing of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$ and $f_{\mathrm{ms}}$ respectively beat together, producing an RF comb in the radio-frequency domain. This RF comb inherits the $[0,-\pi,-\pi,-\pi]$ periodic multilevel phase profile of the H-OFC and exhibits twice the line amplitude compared to the C-DCS.
@@ II Principles
In this section, two DCS architectures for this CH-EDCS concept are studied: one termed C-H-OFC, employing a coherent-harmonic optical frequency comb (H-OFC) realized by TPM and a conventional optical frequency comb (C-OFC); the other termed H-H-DCS, employing two H-OFCs. For clarity, we use a harmonic number of m = 4 and a TPM sequence of $[\pi,0,0,0]$ as illustrative examples.
II.1 C-H-DCS
@@
As illustrated in Figs. 1 (c) and (d), C-H-DCS utilizes a H-OFC whose repetition rate is approximately equal to 4 times of the repetition rate of the C-OFC (or, more generally, any integer multiple m). The pulses of comb 1, after stepping across the first pulse of the H-OFC, subsequently scan its 2nd to 4th pulses, thus generating four centerbursts instead of a single one on the interferogram within the same time period. Additionally, according to the principle of linear optical sampling 35, 36, the expression for the interference intensity can be written as
@@
where $E_{1}(t),E_{2}(t)$ represents the pulses’ electric field envelopes; n denotes the pulse pair number; $\tau$ denotes the time delay induced by the repetition rate difference; and $\phi_{TPM}$ represents the phase difference introduced by the TPM. For each centerburst, the $\phi_{TPM}$ is constant. Consequently, the TPM imposed on the pulses of H-OFC is finally mapped onto the four centerbursts.
Fig. 2: Illustration of the process of the spectral mode enhancement by the harmonic pulse trains in the frequency domain. The j represents the spectral mode index (derived from $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$). The colored linear arrows with indexes of k = 1, 2, 3, 4 represent the mode vectors from the kth-sub-harmonic pulse train at the frequency of $\omega_{j}$, while the blue rotating arrows indicate the temporal phase modulation (TPM), which rotates all the modes of the 1st-sub-harmonic pulse train (k = 1) by $\pi$. In the absence of this TPM, only the modes of $j=4n$ $(n\in\mathbb{Z}^{+})$ will be present, while the remaining modes will undergo destructive interference, consistent with conventional expectation. The black arrows represent the total mode vectors after superposition, which have twice the amplitude of the sub-modes and exhibit a periodic phase profile of $[-\pi,-\pi,-\pi,0]$.
Fig. 3: Concept of H-H-DCS. (a) Time-domain description of the H-H-DCS. The H-OFC 1 and H-OFC 2 emit two pulse trains with repetition rate of $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$ and $4f_{\mathrm{ms}}$ respectively, both with pulse phases of $[\pi,0,0,0]$. Interference of the two pulse trains within a time period of $1/\Delta f_{\mathrm{s}}$ produces four distinct centerbursts: one at the fundamental oscillation frequency (purple), exhibiting a 4-fold amplitude enhancement over conventional dual-comb spectroscopy (C-DCS); and three additional centerbursts at higher-order frequencies (pink), which are suppressed by low-pass filtering in practical implementations. (b) The equivalent frequency-domain description of the H-H-DCS. Two H-OFCs with mode spacing of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$ and $f_{\mathrm{ms}}$ respectively beat together, producing an RF comb in the radio-frequency domain which has a uniform phase profile and exhibits four times the line amplitude compared to the C-DCS. Note that, the uniform phase profile requires the phase profiles of the two H-OFCs to be accurately aligned.
The spectral behavior of the H-OFC and the interferogram with 4 centerbursts can be elucidated through the interference of the harmonic pulses. The original pulse train is decomposed into a collection of four sub-harmonic pulse trains, each with a period of $1/f_{\mathrm{ms}}$ ($1/\Delta f_{\mathrm{s}}$ for the interferogram) and corresponding to a sub-spectrum with mode spacing of $f_{\mathrm{ms}}$. According to the linear property of the Fourier transform, the spectrum of the H-OFC or the interferogram at frequency $\omega_{j}$ is a vector superposition of the 4 sub-modes at that frequency, given by $E_{j}=\sum_{k=1}^{4}a_{0}\exp{(\omega_{j}t+\phi_{k,j})}$, where $a_{0}$ is the amplitude of the sub-modes. Considering the influence of the relative time delay among these harmonic trains and the TPM, the phase spectrum of the kth-sub-spectrum can be expressed as
The first term in the formula, $-\omega_{j}t_{k}$, arises from the contribution of the Fourier time shift, where $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$ is the angular frequency of the $j_{th}$ mode, $t_{k}=(k-1)/4f_{\mathrm{ms}}$ is the time delay of kth-sub-harmonic pulse train relative to the first sub-harmonic pulse train; the second term, $\phi_{TPM,k}$, is the phase shift added to the kth-train through the TPM. The superposition process and results are illustrated in Fig. 2, with vector sum $E_{j}=2a_{0}\exp{(\omega_{j}t+\varphi_{j})}$ where $\varphi_{j}$ is a periodic sequence $[-\pi,-\pi,-\pi,0]$. This indicates that the spectrum of the original pulse train has the same mode spacing but twice the amplitude of the sub-spectrum, and exhibits a periodic multilevel phase spectrum of $[-\pi,-\pi,-\pi,0]$.
Thus, an SNR gain of 2 can be achieved without coherent averaging 4 IGMs (spanning a total time of $4/\Delta f_{\mathrm{s}}$ in the C-DCS). For a more general case, the number 4 in the aforementioned equations corresponds to m, and the corresponding SNR gain is $\sqrt{m}$. Moreover, it is worth noting that if a flat gain is required, an equation in the quadratic function form indicated by the so-called ”Talbot condition” 37 can be used, which is expressed as
$\phi_{n;s,m}=\pi\frac{s}{m}k^{2}(mod~{}2\pi),$ (3)
where s, m are mutually prime integer numbers, and k is the pulse number. Correspondingly, the periodic multilevel phase profile of the total spectrum is $\varphi_{k;p,q}=-\pi\frac{p}{q}j^{2}(mod~{}2\pi),$ where $q=m$, j is the mode number, and $p\neq s$ in general. Whereas, if a reference channel for spectral normalization in DCS configuration is available, the majority of periodic signal can be set.
II.2 H-H-DCS

### v4

# Reducing Averaging Time in Dual-comb Spectroscopy via Phase-Patterned Higher-Repetition-Rate Pulses

@@ Abstract
Dual-comb spectroscopy (DCS) is a powerful Fourier-transform spectroscopic technique that provides high-speed, high-resolution, and broadband measurements without moving parts. However, the high peak power of mode-locked pulses limits the photodetector’s dynamic range, resulting in a low signal-to-noise ratio (SNR) per acquisition. While coherent averaging can improve SNR, it sacrifices temporal resolution and demands stringent system stability. Here, we introduce a novel concept to enhance SNR by using phase-patterned higher-repetition-rate combs. We reinterpret the self-imaging process of comb spectrum from a new perspective on mode interference among sub-pulse trains As a proof-of-concept, we densified two 250-MHz frequency combs to 12.5-MHz mode spacings via phase modulation and performed DCS on an $\mathrm{H^{13}C^{14}N}$ gas cell, and compared the results with an emulated conventional 12.5-MHz DCS, demonstrating a 17-fold increase in mode amplitude. This concept is expected to be combined with ultra-high repetition rate combs, such as microcombs, and thereby deployed in practical applications that typically require spectral sampling spacings from hundreds of MHz to GHz range.
@@ I Introduction
Enabled by the unique properties of optical frequency combs (OFCs) 1, dual-comb spectroscopy (DCS) has emerged over the past decade as a powerful, next-generation Fourier-transform spectroscopy technique, offering an important tool for applications requiring high speed, broad bandwidth, high frequency resolution, and high frequency accuracy, such as in investigations of physical turbulent environments 2, 3, 4, 5, 6 and chemical and biological dynamic processes 7, 8, 9, 10. By employing two coherent OFCs with a slight repetition rate difference ($\Delta f_{\mathrm{rep}}$), and inherently free from mechanical moving parts, DCS can capture an interferogram (IGM), yielding a spectrum with a spectral resolution set by the repetition rate of the combs. The time frame for this capture is as short as $1/\Delta f_{\mathrm{rep}}$, and with large $\Delta f_{\mathrm{rep}}$ configurations, this leads to theoretical temporal resolution reaching microsecond or even nanosecond timescales 11, 12, 13.
@@
However, in typical practical DCS systems, the high peak power of mode-locked pulses often leads to photodetector saturation or nonlinear response, thus restricting the average power of the combs to a low level. Consequently, given the vast number of comb teeth inherent to broadband spectra, the power per comb tooth is typically limited to the nanowatt range 14, ultimately resulting in extremely low SNR per single acquisition, When the comb source power is sufficient, this is referred to as the dynamic range limitation15, which necessitates coherent averaging or multi-period acquisition to enhance the SNR 16, 17, 12. Furthermore, the SNR gain achieved through averaging scales proportionally to the square root of the number of averages. For instance, to improve the SNR by a factor of 10, a hundred averages are needed. Obviously, this process not only significantly compromises the exceptional time-resolution of DCS by several orders of magnitude, but also places greater demands on the long-term mutual coherence and stability of the system, which greatly increases the implementation complexity and cost.
@@
To mitigate this issue, one approach involves employing sequential spectrum acquisition using multiple photodetectors 14 or averaging spectral copies instead of IGMs 18. However, the SNR enhancement offered by these two methods has been limited by experimental complexity or sampling bandwidth constraints, demonstrating averaging-time reduction of less than one order of magnitude. More recently, the utilization of $\sim$ 1-GHz-repetition-rate combs 12, 19, 13, 20 offers higher average power and larger $\Delta f_{\mathrm{rep}}$. Nevertheless, a $\sim$ 1-GHz repetition rate had struck a delicate balance between measurement speed and spectral bin spacing 21. A higher repetition rate inherently implies a lower spectral bin spacing, which limits further increases in repetition rate.
Over the past two decades, a comb densification technique based on the spectral self-imaging effect (also known as the spectral Talbot effect) 22 has been developed 23, 24, 25, 26, 27, 28. This technique applies periodic temporal-phase modulation (TPM) to the comb pulses, thereby inducing evenly spaced ”self images” of the original comb modes and leading to the densification of the comb 29, 30, 31.
In this paper, we model the modulated pulse train as a collection of multiple sub-pulse trains and reinterpret the self-imaging process of comb spectrum from a new perspective on mode interference among these trains. This perspective is analogous to the supermode analysis in harmonic mode-locking32, 33. Subsequently, we intuitively demonstrate how multiple sub-pulse trains synergistically enhance their common modes. This shows that mode energy can be increased using higher-repetition-rate pulses within the same detector dynamic range without reducing spectral bin spacing, thereby providing an effective solution to the SNR limitation in DCS.
Building on this framework, we provide a complete description for two architectures for this special form of DCS. As a proof-of-concept, we densify two close-to-250 MHz frequency combs via periodic TPM and measured the transmission spectrum of a hydrogen cyanide ($\mathrm{H^{13}C^{14}N}$) gas cell at a spectral sampling bin spacing of 12.5 MHz. Compared to an emulated conventional 12.5-MHz DCS, we demonstrate a 17-fold increase in mode amplitude and an SNR gain of up to $\sim$16.7 times per single acquisition in the non-shot-noise-limited regime, the latter corresponding to a 290-fold reduction in averaging time. This average time reduction is expected to reduce to 17-fold in the shot-noise-limited regime. To distinguish it from conventional dual-comb spectroscopy (C-DCS), we term this approach phase-patterned higher-repetition-rate-pulse enhanced dual-comb spectroscopy (PPHE-DCS).
@@ II Principles
In this section, we first present the spectral behavior of phase-patterned higher-repetition-rate pulses from a perspective of mode interference among sub-pulse trains. Then, two DCS architectures for this PPHE-DCS are studied: one employing a conventional optical frequency comb (C-OFC) and a phase-patterned higher-repetition-rate optical frequency comb (PH-OFC), termed C-PH-DCS; the other employing two PH-OFCs, termed PH-PH-DCS. For clarity, we use a pulse-multiplexing number of m = 4 and a TPM sequence of $\pi,0,0,0]$ as illustrative examples.
II.1 Spectrum of Phase-Patterned Higher-Repetition-Rate Pulses
As a example, when phase modulation an OFC with a repetition frequency of $4f_{\mathrm{ms}}$ with a $f_{\mathrm{ms}}$-periodic sequence of $\pi,0,0,0]$, the original pulse train can be decomposed into a collection of four sub-pulse trains, each with a repetition frequency of $f_{\mathrm{ms}}$, as illustrated in Fig. 1(a). The pulses within each sub-pulse train are still fully coherent, thereby each sub-pulse train corresponding to a sub-spectrum with mode spacing of $f_{\mathrm{ms}}$ (Fig. 1(b)). According to the linear property of the Fourier transform, the spectrum of the PH-OFC or the interferogram at frequency $\omega_{j}$ is a vector superposition of the 4 sub-modes at that frequency, given by $E_{j}=\sum_{k=1}^{4}a_{0}\exp{(\omega_{j}t+\phi_{k,j})}$, where $a_{0}$ is the amplitude of the sub-modes. Considering the influence of the relative time delay among these sub-pulse trains and the TPM, the phase spectrum of the kth-sub-spectrum can be expressed as
The first term in the formula, $\omega_{j}t_{k}$, arises from the contribution of the time shift, where $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$ is the angular frequency of the $j_{th}$ mode, $t_{k}=(k-1)/4f_{\mathrm{ms}}$ is the time delay of kth-sub-pulse train relative to the first sub-pulse train; the second term, $\phi_{TPM,k}$, is the phase shift added to the kth-train through the TPM. The superposition process and results are illustrated in Figs. 1(b) and (c), with vector sum $E_{j}=2a_{0}\exp{(\omega_{j}t+\varphi_{j})}$ where $\varphi_{j}$ is a periodic sequence $\pi,\pi,\pi,0]$. This indicates that the spectrum of the phase-patterned higher-repetition-rate pulses train has the same mode spacing but 2-fold the amplitude of the sub-spectrum (which directly corresponds to a 4-fold increase in pulses, i.e., a 4-fold increase in optical power), and exhibits a periodic multilevel phase spectrum of $\pi,\pi,\pi,0]$. This form of spectrum can be viewed as the result of interleaving four independent modesets, within which the mode spacing is $f_{\mathrm{ms}}$ and the modes are perfectly phase-coherent.
In practice, a deviation in the phase modulation level leads to inconsistent interference of modes at different frequencies when all the sub-spectra are superimposed (imaging that the rotation angle of the blue arrow in Fig. 1(c) is not exactly 180 degrees), which in turn causes periodic mode-to-mode-amplitude variations. This also corresponds to the more general case of using an arbitary signal and the resulting spectrum can still be viewed as the result of interleaving four independent modesets.
When the period of the modulation signal is greater than 4, to obtain a consistent constructive interference, one can use either the quadratic function form indicated by the so-called ”Talbot condition” 34 or a binary pseudo- random signal26. The equation of the former is expressed as
$\phi_{n;s,m}=\pi\frac{s}{m}k^{2}(mod~2\pi),$ (2)
where s, m are mutually prime integer numbers, and k is the pulse number. Correspondingly, the periodic multilevel phase profile of the total spectrum is $\varphi_{k;p,q}=-\pi\frac{p}{q}j^{2}(mod~2\pi),$ where $q=m$, j is the mode number, and $p\neq s$ in general.
Fig. 1: Illustration of the process of the spectral mode superposition in phase-patterned sub-pulse trains. (a) Time domain description. (b) The equivalent frequency-domain description. The j represents the spectral mode index (derived from $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$). The $\pi$ phase of the first spectrum originates from the $\pi$ phase shift in the time domain; the phase slopes of the 2nd to 4th spectra are caused by the phase shift of $\omega_{j}t_{k}$ arising from the temporal shift of pulses. (c) Illustration of vector superposition of the modes of the four sub-pulse trains. The colored linear arrows represent the mode from the kth pulse train at $\omega_{j}$, while the blue rotating arrows indicate the temporal $\pi$-phase shift on the k = 1 pusle train, which rotates all its spectral modes by $\pi$. The black arrows represent the total mode vectors after superposition, which have twice the amplitude of the sub-modes and exhibit a periodic phase profile of $\pi,-\pi,-\pi,0]$. In the absence of the $\pi$-shift, only the modes of $j=4n$ $(n\in\mathbb{N})$ will be present, while the remaining modes will undergo destructive interference. Consequently, the mode spacing is consistent with the temporal pulse repetition rate, at $4f_{\mathrm{ms}}$, which aligns with classic OFCs.
Fig. 2: Concepts of conventional DCS (C-DCS) and C-PH-DCS. (a) Time-domain description of C-DCS. The pulse train (red) of comb 1 progressively steps across another pulse train (blue) with a slight repetition rate difference $\Delta f_{\mathrm{s}}$, generating an interferogram with a single centerburst in a period of $1/\Delta f_{\mathrm{s}}$. When the temporal interval between the two pulses is large, effective interference does not occur. (b) The equivalent frequency-domain description of the C-DCS. The beating of two combs produces an RF comb in the radio-frequency domain. (c) Time-domain description of the C-PH-DCS. The pulse of the conventional optical frequency comb (C-OFC) with a pulse repetition rate of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$, steps across the four pulses of the phase-patterned higher-repetition-rate optical frequency comb (PH-OFC), which have phases of $\pi,0,0,0]$ and a pulse repetition rate of $4f_{\mathrm{ms}}$. This generates four centerbursts within an interferogram period of $1/\Delta f_{\mathrm{s}}$, which inherit the $\pi,0,0,0]$ phases. (d) The equivalent frequency-domain description of the C-PH-DCS. Two combs with mode spacing of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$ and $f_{\mathrm{ms}}$ respectively beat together, producing an RF comb in the radio-frequency domain. This RF comb inherits the $\pi,\pi,\pi,0]$ periodic multilevel phase profile of the PH-OFC and exhibits twice the line amplitude compared to the C-DCS.
Fig. 3: Concept of PH-PH-DCS. (a) Time-domain description of the PH-PH-DCS. The PH-OFC 1 and PH-OFC 2 emit two pulse trains with repetition rate of $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$ and $4f_{\mathrm{ms}}$ respectively, both with pulse phases of $\pi,0,0,0]$. Interference of the two pulse trains within a time period of $1/\Delta f_{\mathrm{s}}$ produces four distinct centerbursts: one at the fundamental oscillation frequency (purple), exhibiting a 4-fold amplitude enhancement over conventional dual-comb spectroscopy (C-DCS); and three additional centerbursts at higher-order frequencies (pink), which are suppressed by low-pass filtering in practical implementations. (b) The equivalent frequency-domain description of the PH-PH-DCS. Two PH-OFCs with mode spacing of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$ and $f_{\mathrm{ms}}$ respectively beat together, producing an RF comb in the radio-frequency domain which has a uniform phase profile and exhibits four times the line amplitude compared to the C-DCS. Note that, the uniform phase profile requires the phase profiles of the two PH-OFCs to be accurately aligned.
Fig. 4: Experimental setup. Rb clock: rubidium clock; OFC: optical frequency comb; AWG: arbitrary waveform generator; RFA: RF amplifiers; EOPM/EOIM: electro-optical phase/intensity modulator; OC: optical coupler; PD: photodetector; LPF: low-pass filter; BPF: band-pass filter; EDFA: erbium-doped fiber amplifier; CW: continuous wave laser; DWDM: dense wavelength-division multiplexer.
II.2 C-PH-DCS
@@
As illustrated in Figs. 2(c) and (d), C-PH-DCS utilizes a PH-OFC whose repetition rate is approximately equal to 4 times of the repetition rate of the C-OFC (or, more generally, any integer multiple m). The pulses of comb 1, after stepping across the first pulse of the PH-OFC, subsequently scan its 2nd to 4th pulses. Consequently, four centerbursts are generated within a single period of $1/\Delta f_{\mathrm{s}}$, as opposed to the single one generated in C-DCS (Fig. 2(a)). Additionally, according to the principle of linear optical sampling 35, 36, the expression for the interference intensity can be written as
@@
where $E_{1}(t),E_{2}(t)$ represents the pulses’ electric field envelopes; n denotes the pulse pair number; $\tau$ denotes the time delay induced by the repetition rate difference; and $\phi_{TPM}$ represents the phase difference introduced by the TPM. For each centerburst, the $\phi_{TPM}$ is constant. Consequently, the TPM imposed on the pulses of PH-OFC is finally mapped onto the four centerbursts.
The spectral behavior of the interferogram with four phase-patterned centerbursts can be also deduced from the spectral mode superposition process as illustrated in Fig. 1. Consequently, the C-PH-DCS (Fig. 2(d)) provides a 2-fold amplitude-enhancement in spectral mode compared to the C-DCS (Fig. 2(b)), all while preserving the spectral sampling spacing. For a more general case, the number 2 aforementioned corresponds to $\sqrt{m}$, where m is defined as the multiplexing factor of the pulses.
II.3 PH-PH-DCS